A woman throws her keys straight up at 8.49m/s. Her friend catches them 2.89m higher. How long were the keys in the air
1
Expert's answer
2020-06-08T10:16:08-0400

The equation for the path made by a body thrown upward with initial velocity "v" is


"h=vt-gt^2\/2,"

express time:


"t=\frac{v+\sqrt{v^2-2gh}}{g}=1.27\text{ s}."